

********************************************************************
UARS WEEKLY STATUS REPORT - 24 Dec 2004
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
The SUSIM experiment successfully gathered daily mid (1.1 nm) resolution,
MgII, occultation, solar pointing offset, and continuous monitoring
scans for time periods with UARS telemetry or tape recording.
Thermospheric molecular oxygen profiles have been derived from
the SUSIM solar occultation data for 1991-1997. The results were
presented to the Fall Meeting of the AGU. A copy of the poster
(Powerpoint) may be obtained from:
http://solphys.nrl.navy.mil/Thermospheric_Occultation.ppt
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 11 December 2004 (Orbit 72519, GMT Day 346) through
17 December 2004 (Orbit 72623, GMT Day 352).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 349/1627 72574 POWER OFF
PEM 350/1308 72587 AXIS 1&2 ON
PEM 351/1402 72603 ZEPS/NEPS FULLTIME
The beta angle ranged between the angles of 30.5 degrees to 63.3
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 80 degrees on 23 December 2004 (DOY 358).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.26 to 1.59 degrees C. Additional
battery performance data for each battery on 03 December and 10
December 2004 are provided in Appendix D.
The clock error ranged from 10.2 to -10.6 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 03
PEM = 02
HALOE = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 02
Zero (0) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continued their CMS software acceptance testing this week.
This testing will continue into the new year.
The UARS FOT would like to extend their best wishes for this holiday
season to all those who helped make this another successful year for
UARS.
MEETINGS
-None this report period
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 12 January 2005 (DOY 012).
This maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 17 December 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
72520 346/0850 04-344 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
72537 347/1331 04-345 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
72574 349/1250 04-346 HALOE POWER OFF (TBAT)
72587 350/1253 04-347 PEM AXIS 1&2 ON
72591 350/0850 04-348 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
72603 351/1253 04-349 PEM ZEPS/NEPS TO FULLTIME
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 17 December 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 17 December 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
72520 346/0217 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
72537 347/0617 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
72591 350/1951 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 17 December 2004
10 December 2004 - GMT Day 345 Beta = 30.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -2.73/-3.68 +00.0/+00.0 07.0 0.00 0.0
2 44.8/ -05.6 +6.44/+5.47 +28.0/-13.6 25.9 1.03 20.7
3 -50.4/-190.4 +5.15/+4.51 +26.8/-13.2 25.8 1.02 20.2
17 December 2004 - GMT Day 352 Beta = 61.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.99/-4.94 +00.0/+00.0 07.0 0.00 0.0
2 -11.2/ -56.0 +4.50/+3.23 +21.6/-13.6 26.9 1.05 17.0
3 -33.6/-140.0 +2.91/+1.97 +20.0/-12.8 26.7 1.04 16.3
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period
APPENDIX E
UARS ANOMALY REPORTS
Week ending 17 December 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-090 FOT 72523 346/0705 30832 Generic Late Acquisition
04-091 FOT 72530 346/1854 37847 No UPD/GCMR Capability
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-090 FOT See List Dec 04 30832 Generic Late Acquisition
04-091 FOT 72530 346/1854 37847 No UPD/GCMR Capability